# Rethinking CLIP-based Video Learners in Cross-Domain Open-Vocabulary Action Recognition

### v1

@@ 5.3 Conclusion
This work concentrates on a valuable but underexplored task, namely cross-domain open-vocabulary action recognition. By contributing the XOV-Action benchmark, we provided a comprehensive way to evaluate models across various types of domain gaps and different action categories, and our evaluation of state-of-the-art CLIP-based video learners revealed the challenging nature of this valuable task. To address this task, we proposed a novel scene-aware video-text alignment method, aiming to tackle the domain gap by learning scene-agnostic video representations. Extensive experimental results demonstrated the effective of our method.
@@
The cross-domain open-vocabulary action recognition task presents numerous challenges that remain to be addressed, such as cross-category generalization in cross-domain settings. Despite our efforts, we focus only on resolving scene bias in our current methodology. However, our proposed benchmark provides a comprehensive way for evaluation. We believe that our work will serve as a catalyst for further advancement, and we hope it will inspire future innovative solutions for this field.

### v2

@@ 6 Conclusion
This work concentrates on a valuable but underexplored task, namely cross-domain open-vocabulary action recognition. By contributing the XOV-Action benchmark, we provided a comprehensive way to evaluate models across various types of domain gaps and different action categories, and our evaluation of five state-of-the-art CLIP-based video learners revealed the challenging nature of this valuable task. To address this task, we proposed a novel Scene-Aware video-Text Alignment (SATA) method, aiming to mitigate the scene bias when fitting training videos. Quantitative comparison on XOV-Action showed that our SATA can effectively improve the closed-set action recognition performance across domains while generally maintaining the open-set performance, and extensive analysis demonstrated the effectiveness of our method.
@@
This work focuses on resolving the scene bias in current methodology, which mainly targets the cross-domain generalization for closed-set categories (i.e., Challenge I discussed in Section 3). Despite our efforts, the task of cross-domain open-vocabulary action recognition still presents numerous challenges, including improving open-set recognition across domains and developing a unified model for all cases (Challenges II and III in Section 3). Fortunately, our proposed XOV-Action benchmark provides a comprehensive way for evaluation and analysis. We believe that our work will serve as a catalyst for further advancement, and we hope it will inspire future innovative solutions for this field.